OpenStudy (anonymous):

what is y=2/7x-4 written in standard form?

OpenStudy (anonymous):

Standard form is ax + by = c y = 2/7x - 4 y + 4 = 2/7x 4 = 2/7x - y Multiply everything by 7 to get rid of fraction. 28 = 2x - 7y Answer: 2x - 7y = 28